Jessica Banks, Amanda Parkes

Cambridge, MA USA

Curiously Strong (2004)

Altoids boxes, solenoids, acrylic, copper

A string of actuated Altoids(TM) boxes is mechanically controlled so that the lids open and close in sequence. Because the piece is long and flexible, it can be configured to take advantage of (or interfere with) the natural organization of a room and can be experienced by many people at the same time - both visually and audibly. The falling of the lids and clicking of the solenoids creates a rhythmic, moving soundscape, enveloping the space like a Sisyphean loop of dominoes. Curiously Strong is funded in part by a Director’s Grant from the Council for the Arts at MIT.

Joëlle Bitton, Stefan Agamanolis, Matthew Karau

Dublin, Ireland

http:://www.medialabeurope.org/hc/raw

RAW (2003)

Sound+still pictures+ISIS

RAW is a system combining a tool and a process for capturing and conveying audiovisual impressions of everyday life. The tool itself incorporates a digital camera and a binaural audio recording device that captures the minute of sound before and after a picture is taken. In August 2003, the tool was taken to Mali in Africa and given to 23 people who expressed themselves on their daily lives, recording sound and taking pictures, without editorial direction. The installation presents their content, without alteration, and allows the audience to immerse itself in a collection of moments of humanity. The project aims to enable a relationship between the user of the tool and an audience in a different place or time with an absolute minimum of editorial mediation by a third party. It is also addressing the questions of the editing process and mediation in audiovisual narrative experience.

Nell Breyer, jackbackrack

Cambridge, MA USA

http:://www.media.mit.edu/~nbreyer

Walking Wall (2004)

Projection, live video processing with Gooze

The Walking Wall reflects narrow impressions, views, handmade marks and footprints of each visitor’s idiosyncratic dance.

Andrew Zoz Brooks

Cambridge, MA USA

http:://mit.edu/zoz/www/
http:://mit.edu/mcanulty/www/

Funkenschnorkel (2004)

Mixed media construction; audio output

Funkenschnorkel (corrupted German: “broadcasting snorkel”) explores empowering the individual to take control of the audio environment of medium to large scale interior and exterior spaces, in order to manipulate the perceptions of other occupants of the space and to explore the combination of sound and space without prior permission. The input combination of digital audio mass storage and microphone with spatial effects provides the wearer with a large repertoire and enhanced vocal capacity. For this static installation, activity of the Funkenschnorkel wearer is simulated using music designed by Dan McAnulty to respond to the motions of the audience. Dr Thomas von Wiegand and Blake Brasher contributed to the onboard and sensor electronics respectively.

Ben Dalton

Cambridge, MA USA

http:://www.media.mit.edu/~bcd

Rabbit Field (2004)

Satin, flannel, PC fans, electronics, Java

Electronically connected inflatable shapes. Squeeze softly to deflate. Thanks go to Carlos A. Rocha for the development of rabbitNet.

Jeana Frost

Cambridge, MA USA

Cables and Wires (2004)

Mixed method woodcuts

For these images, original digital photographs are first color separated. These images are then laser cut into wood. The woodblocks are further cut and manipulated by hand and then printed on paper.

jackbackrack

Cambridge, MA USA

http:://www.jbot.org

Intimacy Machine (2004)

Interactive video sculpture with one pc and two each of computer video cameras, lcd panels, two way mirrors, and boxes.

The intimacy machine is a computer based reciprocal peepshow. It mediates intimacy allowing people to overcome their normal social boundaries. In particular it provides an indirect mechanism for people standing in close proximity to each other to stare at each other directly in the eyes, a feat that otherwise proves tremendously difficult for humans. Whereas telesex offers a way for people far away to feel close, the intimacy machine makes it possible for people that are close to feel distant. It also offers a facility for people to watch themselves watching. Finally, in other moments the machine reveals an ever changing intimate small world generated from macro images of self. The artist would like to thank Dan Paluska, Mindy Zarem, Jeff Weber, Jessica Banks, Fran Trainor, Randall Heath, and Nell Breyer for their help in construction and otherwise.

Brian Knep

Boston, MA USA

http:://www.blep.com

Architectural Drip (2004)

Computer, projector, camera

An organic pattern grows and flows along one of the gallery pillars. The piece inhabits this structural element and becomes a living architectural detail.

Nick Knouf

Cambridge, MA USA

http:://www.zeitkunst.org

44#444 (2004)

Dumb terminals, point-of-sale receipt printer, custom software.

44#444 is an interactive installation that engages participants in a reexamination of modern methods of communication, namely cell-phone text messaging, by twisting representations into new forms. Funded (in part) by a Director’s Grant from the Council for the Arts at MIT.

Jeff Lieberman

Cambridge, MA USA

http:://bea.st

Slink (2004)

Aluminum, Corroded Steel, Acrylic, Electronics, Custom Voice Coil with Flexure Mounts, 1980 LEDs, Extension Spring.

An experiment with matched mechanical, electrical, and visual resonances, using a spring and light. Many thanks to: Jack Holloway (hardware), Josh Glazer (software), Neil Jenkins (materials), and to jackbackrack, James Bales, Zoz Brooks, Arthur Ganson, Eric Gunther, Jessica Hinel, Dan Lovell, Kevin McKormick, Dinsha Mistree, Cha-Ling O’Connell, Dan Paluska, Derek van Beever, Dick Whitney, for soldering time, useful conversations, and inspirado.

Andrew Neumann

Boston, MA USA

http:://www.bitforms.com/artist neumann.html

Quartet (2003)

Wood, LCD screens, cameras, electronics

Nancy Paterson

Toronto, Canada

http:://www.vacuumwoman.com

STOCK MARKET SKIRT (2000)

Internet-Interactive

A cyberfeminist fashion statement.

James Patten

Cambridge, MA USA

http:://www.media.mit.edu/~jpatten

Three rotations (2004)

Steel, aluminum, wood, rubber, foam, motors, electronics, computers

Three rotations explores three different relationships between objects and sound. The viewer is encouraged to touch each one, and move around them to hear the sound space created by the three together.

Hayes Solos Raffle, Dan Maynes-Aminzade

Cambridge, MA USA

web.media.mit.edu/~hayes/mas863/urinecontrol.html

You’re In Control (Urine Control) (2002)

Urinal, electronics, PIC microcontroller, PC gaming equipment running custom video game

Play while you pee! With You’re In Control, the urinal knows your every move. Strap on the Game Controller and have a squirt at the flying hamsters. Hit them all for the high score. Improve your aim and improve your game.

Ready to “flush” those hamsters? See if you can “handle” our hidden bonus game!

Ramesh Raskar, Amit Agrawal

Cambridge, MA USA

http:://raskar.com

Camera Non-Photo (2004)

Multi Flash Camera + Projector + PC

Photorealism in visual arts is passé. An artist creates a rendering that is not-a-photo. What about a camera that produces not-a-photo? A camera with strategically positioned flashes cast shadows along silhouettes along different directions in the scene. The multi-flash camera detects shadows and ignores colors to capture the shape boundaries and depicts just the important lines and edges, so the observer can use the perceptual cues to fill in the rest of the image. The result is a computational camera generating A-Ha’s “take on me” line drawing-style animation video of live interactions of participants.

Dan Roe

Cambridge, MA USA

http:://www.danroe.net

Specimen (2004)

Solar engines, steel

Specimen examines the conflict between our desire to understand the natural world on our terms, and the natural world’s desire exist on its terms. But specimen is nothing but steel, glass, and a few assorted electronic components. Ironically, it is our own curiosity that is on display.

Fran Trainor

Somerville, MA USA

Nine Cell Studies (2004)

Epson 2200 prints

A series of small studies exploring the relationship between cultural and biological production.

William Tremblay

Allston, MA USA

Bionic Log (2003)

Wood, steel, latex, electronics

Bionic Log is a pneumatic robotic sculpture built around a section of tree trunk. Inside the log is a bank of pneumatic valves driven by a simple microcontroller program. These valves permit the flow of compressed air to the actuators in the limbs of the robot, which contract in a manner very similar to animal muscles. Although the control mechanism is simple, the resulting motions are very recognizable as human gestures, albeit the gestures of a confused or wounded person. Bionic Log is an extrapolated collision of the conflicting human imperatives of expediency and sentimentality, at once supporting and disproving the comforting notion that technology can solve any problem.